Circular No. 5995

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

CATALOGUE OF COMETARY ORBITS 1994 A new edition—the ninth—of the Catalogue of Cometary Orbits is

now available. The 103 pages contain orbital elements (in the J2000.0 system) for 1417 cometary apparitions through April 1994. The price, postage included, is $20.00 ($30.00 by airmail outside North America). The main part of the catalogue and the table of ‘original’ and ‘future’ 1/a values for the 298 long-period orbits of the highest quality can be supplied by e-mail for $50.00; they are also available on an MS-DOS diskette (5.25inch or 3.5-inch) for $110.00 ($120.00 for airmail delivery), this including a facility for extracting individual orbits and computing ephemerides. Checks should be made out to ‘Central Bureau for Astronomical Telegrams’ and mailed to the postal address given above; subscribers to these Circulars can have their accounts debited by the appropriate amount.

COMET MUELLER (1993p)

Ephemeris extension from orbital elements on MPC 23106:

1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 28 9^{h}11.^{m}18 −23^{o}28.^{′}0 1.118 1.430 84.^{o}1 44.^{o}8 12.8
June 2 9 33.39 −18 55.8 1.224 1.489 82.8 42.5 13.2
7 9 51.75 −15 07.1 1.342 1.548 80.8 40.3 13.5
12 10 07.29 −11 57.3 1.467 1.608 78.4 38.2 13.9
17 10 20.73 − 9 20.7 1.599 1.669 75.7 36.1 14.2
22 10 32.58 − 7 11.5 1.733 1.730 72.8 34.1 14.6
27 10 43.22 − 5 24.6 1.870 1.792 69.7 32.1 14.9
July 2 10 52.90 − 3 56.1 2.007 1.854 66.5 30.2 15.2
7 11 01.85 − 2 42.5 2.145 1.915 63.2 28.3 15.5
12 11 10.19 − 1 41.3 2.280 1.977 59.9 26.4 15.8
17 11 18.05 − 0 50.3 2.415 2.039 56.5 24.6 16.0

P. Camilleri, Cobram, Victoria, Australia, reports that by May 11 the comet’s condensation was no longer visible in his 0.20-m reflector, the comet looking like a nebulous streak some 1′ × 3′ in size and similar to the illfated comet Bradfield 1992 XIII in late May 1992. Camilleri’s total visual magnitude estimates indicate that the comet is much fainter than expected: May 5.42 UT, 10.0; 11.38, 11.0; 12.38, 11.2.

1994 May 25 Daniel W. E. Green